# The Bayesian Separation Principle for Data-driven Control

### v1

@@ Abstract
This paper investigates the existence of a separation principle between model identification and control design in the context of model predictive control. First, we elucidate that the separation principle holds asymptotically in the number of data in a Fisherian setting, and universally in a Bayesian setting. Then, by formulating model predictive control within a Gaussian regression framework, we describe how the Bayesian separation principle can be used to derive explicit, uncertainty-aware expressions for the control cost and optimal input sequence, thereby bridging direct and indirect data-driven approaches.
Index Terms:
@@ I Introduction
In recent years, several model-based control design problems have been translated into a direct data-driven control framework, including optimal linear-quadratic control [16, 11] and model predictive control (MPC) [13, 6]. However, only few works have discussed the potential advantages of direct approaches over classical indirect methods, e.g. [19, 27]. Notably, [18] suggests that the primary advantage of direct methods may stem from the lack of a separation principle between model identification and control design.
The main objective of this paper is to shed light on the (non)existence of such separation principle by focusing on model predictive control. Specifically, we argue that, in a Fisherian setting, the separation principle holds asymptotically as the data size grows, while, in a Bayesian setting, it holds regardless of the data size. We then delve deeper into Bayesian model predictive control, framing the problem within a Gaussian regression framework and deriving explicit expressions for the control cost and optimal input sequence in some relevant scenarios. Through theoretical insights and numerical examples, we illustrate how controllers based on the Bayesian separation principle can outperform traditional design techniques in terms of control performance. Additionally, throughout the paper, we clarify some frequently used yet often ambiguous terminology, such as the notions of control bias, separation principle for identification and control, direct/indirect and model-free approaches.
Finally, we point out that our work is the continuation of [41, 9] and is similar in spirit to [12] where only linear systems are treated. Moreover, some papers, e.g. [25], have combined Gaussian processes with MPC to enhance control safety, though in a way different from our approach.
II System identification and the classical approach
@@
The aim of system identification is to obtain a mathematical model of a dynamic system starting from input-output measurements. An important aspect is that model construction has often be seen as a problem of relevance. The primary interest might be to predict future data over certain operating conditions. A model might predict well one-step-ahead, e.g. simply setting the prediction to the last measured output, but very poorly at large horizons whose limit case (infinite horizon) requires the system to be simulated without using any past output [43]. Alternatively, one could ask for a model that provides a good description of the system just over frequencies of interest for the control. For MPC this could be again translated into prediction capability over a specific range of horizons, possibly using only a restricted class of inputs.
@@
In the classical approach to system identification described in popular books like [29, 46], the simplest scenario involves a single model structure $\mathcal{M}$. Examples are FIR and ARX in the linear setting or their nonlinear extensions like the NARX model which links inputs $u(t)$ (assumed deterministic) to outputs $y(t)$ as follows
$y(t)\!=\!f(y(t-1),y(t-2),\ldots,u(t),u(t-1),\ldots;\theta)+e(t)$ (1)
@@
with $e(t)$ being white noise. The model depends on an unknown (finite-dimensional) parameter vector $\theta$. The resulting parametrized structure $\mathcal{M}(\theta)$ has to be estimated from $N$ input-output data $\{u(t),y(t)\}$ collected during an experiment and contained in a set denoted by $\mathcal{D}$. Assume for the time being $y(t)$ scalar with $\hat{y}(t|t-1;\theta)$ to indicate its one-step-ahead predictor induced by the model. The latter depends on the outputs only up to instant $t-1$ and corresponds to $f(y(t-1),\ldots,u(t),\ldots;\theta)$ starting from (1). Then, the unknown parameter vector can be obtained via minimization of the prediction error (PEM):
$\hat{\theta}=\arg\min_{\theta}\sum_{t=1}^{N}\big{(}y(t+1)-\hat{y}(t+1|t;\theta)\big{)}^{2}.$ (2)
@@
The performance of this estimator in reconstructing $\theta$ can be e.g. measured in terms of mean squared error (MSE). Using $\mathbb{E}$ to indicate the expectation operator, it consists of the following function of $\theta$:
@@
where $\theta_{i}$ is the $i$th component of the unknown vector. The last equality points out the decomposition of the error into its variance and bias component. Estimators such that
@@
have zero bias component and are therefore called unbiased.
@@
PEM has a long and fruitful history in system identification. Under Gaussian noises and identifiability assumptions, it enjoys some optimality properties. Its link with maximum likelihood (ML) in fact guarantees that it cannot be outperformed by any other unbiased estimator as the data set size grows to infinity (the Cramer-Rao bound will be reached) [10]. Such property also transfers to the estimate of any function of $\theta$ just replacing the unknown parameter vector with its PEM estimate in virtue of the ML invariance principle [50]. We will come back to talk extensively around this point later also in the context of MPC.
@@
Note that the classical approach and its properties are Fisherian in nature: one postulates that there exists a true deterministic system $\mathcal{S}$ that has generated the data and tries to learn it using only the experimental data. Asymptotic optimality of the estimation procedure then holds assuming that the postulated structure $\mathcal{M}$ contains the true system $\mathcal{S}$. However, this scenario is often too simplified since also model complexity is typically unknown, in particular the dimension of $\theta$ which is related to system memory (the number of past input-output data $f$ in (1) depends on). More than one structure needs to be introduced and criteria such as AIC or BIC are used to determine the most appropriate one. They try to find a good trade-off between the two MSE components reported in (3) since, as model complexity augments, the variance increases while the bias decreases. Assuming known (for the sake of simplicity) the variance $\sigma^{2}$ of the stochastic noise $e(t)$ which influences the system, the final model returned by the identification procedure is $\mathcal{M}(\hat{\theta})$ where now
$\hat{\theta}=\arg\min_{\theta}\sum_{t=1}^{N}\frac{\big{(}y(t+1)-\hat{y}(t+1|t;\theta)\big{)}^{2}}{\sigma^{2}}+J(N)\dim(\theta)$
@@
with optimization performed also w.r.t. the dimension of $\theta$. Interestingly, the choice of the penalty term $J$ present in (II) is also a problem of relevance at least in terms of asymptotic properties. AIC, obtained using $J(N)=2$, enjoys minimax properties that may lead to advantages in prediction. However, it overestimates the model order with finite probability. BIC, defined by $J(N)=\log(N)$, is not minimax but asymptotically will return the correct order assuming that this latter exists (parametric scenario) [49]. For an in-depth discussion on the properties of estimators related to model order selection see also [28].
III MPC: direct and indirect approach
@@
Consider now MPC where, in the first instance, the system $\mathcal{S}$ is assumed perfectly known. In addition, the inputs $u$ are deterministic and manipulable. After observing $y^{t-1}$, i.e. the system outputs up to instant $t-1$, the problem is to calculate future inputs $u(t),u(t+1),\ldots$ that allow the system to follow a reference output $y_{\text{ref}}(t)$. Knowledge of $\mathcal{S}$ allows us to define the optimal $\ell$-step ahead predictors $\hat{y}(t+\ell|t)$ by setting them to the mean of future outputs. In this way the predictors minimize the mean squared error on output prediction, and the ideal control cost function to be minimized w.r.t. $u$ is formulated as:
$J_{u}$ $=\sum_{\ell=0}^{T-1}\|y_{\text{ref}}(t+\ell)-\hat{y}(t+\ell)\|_{Q_{\ell}}^{2}$
$+\sum_{\ell=0}^{T-1}\|u_{\text{ref}}(t+\ell)-u(t+\ell))\|_{R_{\ell}}^{2}$ (5)
@@
where here, and in what follows, $\hat{y}(t+\ell)$ is used in place of $\hat{y}(t+\ell|t)$ to simplify notation. In (III) the outputs can now be also vectors, $Q_{\ell}$ and $R_{\ell}$ are positive semidefinite matrices, $\|x\|_{M}^{2}:=x^{\top}Mx$ for any positive semidefinite matrix $M$ and $u_{\text{ref}}$ are reference inputs that complete the definition of the penalty on the energy required by the control.
@@
In a more realistic MPC scenario $\mathcal{S}$ is not known and system identification can enter the scene first. Following our previous discussion, we can introduce a parametric model structure $\mathcal{M}(\theta)$ to capture $\mathcal{S}$. All the induced predictors become $\hat{y}(t+\ell;\theta)$, function of the unknown vector $\theta$. In turn, this leads to the parametrized control cost $J(u;\theta)$. In this setting the simplest version of the so-called indirect approach to data-driven control, of which the classical identification mentioned above is just an example, first determines $\theta$ from the input-output data $\mathcal{D}$ collected in a previous experiment. This can be done through PEM, i.e. minimizing the one-step-ahead prediction error as described in (2) (or solving the more complex (II) in the case of unknown complexity). Then, the objective $J(u;\theta)$ is minimized by replacing $\theta$ with the realization of its estimator $\hat{\theta}$.
@@
There are some recurring criticisms of the indirect approach that have emerged in recent years. The first one is that bias is introduced in the control objective [17]. However we note that such issue will depend on the indirect approach adopted. For instance, classical system identification is parametric in nature but one could also consider different non parametric and/or Bayesian settings as we will do later on. Furthermore a careful definition of bias is needed. To this regard,
@@
in the Fisherian context where $\theta$ is deterministic, we say that the indirect approach to control is unbiased if, for any input $u$, one has
$\mathbb{E}[J(u;\hat{\theta})]=J(u;\theta)\quad\forall\theta,$
with expectation taken w.r.t. the distribution of the estimator $\hat{\theta}$.
@@
If the model structure $\mathcal{M}$ contains $\mathcal{S}$ (so that there exists $\theta$ leading to the optimal predictors) this guarantees that the mean of the cost corresponds to the ideal objective (III).
@@
Adopting the classical PEM approach, the unbiasedness property hardly holds because $J$ is a nonlinear function of $\theta$. On the other hand, when e.g. the identification data $\mathcal{D}$ are nonlinearly related to the model parameters, $\hat{\theta}$ is rarely an unbiased estimator of $\theta$ itself [^1] This is by no means in itself a negative thing. We are here concerned with unbiased costs but there exist estimators that introduce a small bias to greatly decrease variance, hence leading to a favourable mean squared error. Stein also derived in the linear regression setting biased estimators that outperforms least squares for any possible value of $\theta$, e.g. see [20] and [36, Chapter 1].. For this reason one could try to remove this distortion exploiting the experimental observations in a different way. Indeed, the idea that permeates the direct approach is that the information contained in the data should be maximally exploited by focusing directly on the control cost. So,
@@
the problem of relevance related to the direct approach to data-driven control is to build directly from input-output data an estimator of (ideal) control objectives like (III) having desired statistical properties (like unbiasedness).
@@
IV Separation principle for identification and control

### v2

@@ Abstract
In this paper we investigate the existence of a separation principle between model identification and control design in the context of model predictive control. First, we clarify that such a separation principle holds asymptotically in the number of data in a Fisherian context, and show that it holds universally, i.e. regardless of the data size, in a Bayesian context. Then, by formulating model predictive control within a Gaussian regression framework, we describe how the Bayesian separation principle can be used to derive computable, uncertainty-aware expressions for the control cost and optimal input sequence, thereby bridging direct and indirect data-driven approaches. Numerical results in both linear and nonlinear scenarios illustrate that the proposed approach outperform nominal methods that neglect uncertainty, highlighting the advantages of incorporating uncertainty in the control design process.
keywords:
@@ 1 Introduction
In recent years, several model-based control design problems have been translated into a direct data-driven control framework, including optimal linear-quadratic control [12, 8] and Model Predictive Control (MPC) [10, 4]. However, only few works have discussed the potential advantages of direct approaches over classical indirect methods, see e.g. [15, 21, 26]. Notably, [14] suggests that the advantage of direct methods may stem from the lack of a separation principle between model identification and control design.
The primary objective of this paper is to shed light on the existence/non-existence of such a separation principle in the context of MPC. Of course, an answer to this question depends on the chosen probabilistic modeling framework, which defines how uncertainty in the system parameters is represented and handled. In the first part of the paper, we investigate the classic Fisherian setting [23], where system dynamics are governed by deterministic yet unknown parameters and, for known model structure, uncertainty stems from measurement noise. We then consider the Bayesian setting [33], where these parameters are treated as random variables and uncertainty is modeled using prior and posterior distributions.
After establishing a precise definition of separation principle, we highlight an important yet often underappreciated result in the Fisherian context. Namely, assuming e.g. Gaussian noises and quadratic loss functions, it is known that classic prediction error methods yield the optimal parameter estimate asymptotically in the number of data, provided that the class of parametrized models contains the true system. Such estimate in turn yields the optimal control, by virtue of the maximum-likelihood invariance principle [46]. This means that a separation principle between identification and control holds asymptotically in the number of data in the Fisherian setting. For finite data, instead, the separation principle may not hold. In contrast, the Bayesian framework inherently incorporates uncertainty about the system parameters through the posterior distribution, which encodes all available information about the model. We show that this posterior can then be leveraged to directly account for uncertainty in the MPC cost. The optimal control is then directly computed by minimizing this modified cost. Consequently, the separation principle holds universally in the Bayesian case, regardless of the amount of data.
In the second part of the paper, we delve deeper into Bayesian model predictive control, framing the problem within a Gaussian regression framework [37]. We connect it to approaches called model-free in recent control literature. In fact, even when nonlinear systems must be treated, it allows the identification step to be carried out starting from large hypothesis spaces, circumventing ill-posedness simply by including information on the input-output map regularity. This framework permits also to derive explicit expressions for the control cost in some relevant scenarios, treating the nonlinear and linear cases separately. For the nonlinear case, we show that closed-form expressions of the MPC cost can be easily derived for NFIR systems and provide a method for obtaining approximate, computationally tractable expressions for NARX systems. Instead, in the linear case, we leverage the known structure of $k$-step ahead predictors to obtain computable expressions for both the control cost and optimal input sequence. We test our proposed approach, termed Bayesian Separation Principle (BSP), in both a simple, analytically tractable linear scenario and a nonlinear case study. In both scenarios, BSP outperforms the nominal approach, which defines the MPC cost using mean parameter values and neglect prediction uncertainty.
Additionally, throughout the paper, we not only clarify the definition of separation principle between identification and control, but also provide precise definitions of other frequently used yet often ambiguous terminology, such as the notions of control bias and of the aforementioned model-free control. By offering clear definitions and contextual explanations, we aim to resolve any ambiguity and foster a more accurate understanding of these concepts and their relevance in data-driven control.
To conclude, we point out that there are existing works that share similarities with ours. In particular, our work can be seen as a continuation of [38], where Bayesian kernel-based methods have been exploited to design optimal closed-loop controllers for linear systems (for a more comprehensive overview of this and related topics, see [6]). More recently, [9] framed MPC design within a Bayesian framework by focusing on the linear setting and deriving closed-form expressions for optimal inputs that account for prediction uncertainty. While similar in spirit, our work deals with the more general nonlinear setting and, even in the linear case, the control expressions we derive differ from those in [9], as detailed in the numerical section of the paper. Finally, [19, 20] have combined Gaussian processes with MPC to enhance control safety. Differently from our approach, in these papers the authors consider an hybrid modeling framework which combines a nominal system description with an additive Gaussian process term accommodating uncertainty.
2 System identification and the classical approach
@@
The aim of system identification is to obtain a mathematical model of a dynamic system starting from input-output measurements. An important aspect is that model construction has often to be seen as a problem of relevance. The primary interest might be to predict future data over certain operating conditions. A model might predict well one-step-ahead, e.g. simply setting the prediction to the last measured output, but very poorly at large horizons whose limit case (infinite horizon) requires the system to be simulated without using any past output [40]. Alternatively, one could ask for a model that provides a good description of the system just over frequencies of interest for the control. For MPC this could be again translated into prediction capability over a specific range of horizons, possibly using only a restricted class of inputs.
@@
In the classical approach to system identification described in popular books like [23, 42], the simplest scenario involves a single model structure $\mathcal{M}$. Examples are FIR and ARX in the linear setting or their nonlinear extensions like the NARX model which links inputs $u(t)$ (always assumed deterministic in what follows) to outputs $y(t)$ as follows
$y(t)\!=\!f(y(t-1),\ldots,y(t-m),u(t),\ldots,u(t-m);\theta)+e(t)$ (1)
@@
where $e(t)$ is white noise while the integer $m$ defines the system memory. The model depends on an unknown (finite-dimensional) parameter vector $\theta$. The resulting parametrized structure $\mathcal{M}(\theta)$ has to be estimated from $N$ input-output data $\{u(t),y(t)\}_{t=1}^{N}$ collected during an experiment and contained in a set denoted by $\mathcal{D}$. Assume for the time being $y(t)$ scalar with $\hat{y}(t|t-1;\theta)$ to indicate its one-step-ahead predictor induced by the model. This predictor depends on all the outputs up to instant t1t-1 and corresponds to $f(y(t-1),\ldots,u(t),\ldots;\theta)$ when the NARX model (1) is adopted. The unknown parameter vector can then be estimated via the prediction error minimization (PEM) approach:
$\hat{\theta}=\arg\min_{\theta}\sum_{t=1}^{N}\big(y(t)-\hat{y}(t|t-1;\theta)\big)^{2}.$ (2)
@@
The performance of this estimator can be measured in terms of the mean squared error (MSE). The MSE is computed before seeing the identification data, with $\hat{\theta}$ therefore being interpreted as a random vector because of the dependence on the stochastic outputs $y(t)$. Using $\mathbb{E}$ to indicate the expectation operator, it is given by the following function of $\theta$:
@@
where $\theta_{i}$ is the $i$th component of $\theta$. The last equality points out the error decomposition into its variance and bias component. If
@@
the estimator has null bias component and is therefore called unbiased.
@@
PEM has a long and fruitful history in system identification. Under Gaussian noises and identifiability assumptions, the PEM estimator enjoys some optimality properties. Its link with maximum likelihood (ML) in fact guarantees that it cannot be outperformed by any other unbiased estimator as the data set size grows to infinity: the Cramer-Rao bound will be reached making PEM variance as small as possible [7]. Such property also transfers to the estimate of any function of $\theta$ just replacing the unknown parameter vector with its PEM estimate by virtue of the ML invariance principle [46]. We will come back to talk extensively around this point later in the MPC context.
@@
The classical approach and its properties are Fisherian in nature: one postulates that there exists a true deterministic system $\mathcal{S}$ that has generated the data and tries to learn it using only the experimental data. Asymptotic optimality of the estimation procedure then holds assuming that the postulated structure $\mathcal{M}$ contains the true system $\mathcal{S}$. However, this scenario is often too simplified. Also model complexity is typically unknown, in particular the dimension of $\theta$ which is related to the system memory $m$ in (1). More than one structure needs to be introduced and criteria such as AIC or BIC are used to determine the most appropriate one. Such criteria try to find a good trade-off between the two MSE components reported in (3). In fact,
more parameters provide more model flexibility and allow a better fit of the identification data, decreasing the bias. But more parameters also allow adjustments to the noise in the estimation data, augmenting the variance term. The improved fit due to such adjustments has no value and can decrease model prediction capability [^1] While this classical principle of parsimony suggests that the model should describe the experimental data without being too complex, there are however some overparametrized architectures, like deep networks, that only emphasize fitting data and may still perform well on future (unseen) data. For a recent survey on such phenomenon the reader may refer to [29]..
[…]
$\hat{\theta}=\arg\min_{\theta}\sum_{t=1}^{N}\frac{\big(y(t)-\hat{y}(t|t-1;\theta)\big)^{2}}{\sigma^{2}}+J(N)\dim(\theta)$
@@
with optimization performed also w.r.t. the dimension of $\theta$. Interestingly, the choice of the penalty term $J$ in (2) is also a problem of relevance at least in terms of asymptotic properties [45]. AIC, obtained using $J(N)=2$, enjoys minimax properties described in [44] that may lead to advantages in prediction. However, it overestimates the model order with finite probability. BIC, defined by $J(N)=\log(N)$, is not minimax but asymptotically will return the correct order assuming that this latter exists (parametric scenario). For an in-depth discussion on the properties of estimators related to model order selection see also [22].
3 MPC: direct and indirect approach
The arguments developed in this paper concern two distinct phases of an experiment involving a dynamic system: identification, described in the previous section, and control, introduced below. For what regards notation, in what follows it is convenient to use $\tau$ as the starting time of the control phase and assume that the identification data set $\mathcal{D}$ contains input-output pairs observed before $\tau$. For a generic time instant $t$, we also use $y^{t-1}$ to denote the system outputs up to t1t-1. When used as an argument of the NARX system $f$ in (1) this has however to be understood as the vector containing only the previous $m$ outputs. Similar considerations apply to the notation $u^{t}$.
@@ 3 MPC: direct and indirect approach
Consider now MPC where, in the first instance, the system $\mathcal{S}$ is assumed perfectly known. In addition, the deterministic input $u$ is manipulable. After observing $y^{\tau-1}$, the problem is to calculate future inputs $u(\tau),u(\tau+1),\ldots$ that allow the future outputs $y(\tau),y(\tau+1),\ldots$ to follow a reference $y_{\text{ref}}(\tau),y_{\text{ref}}(\tau+1),\ldots$. Knowledge of $\mathcal{S}$ allows us to define the optimal $\ell$-step ahead predictors $\hat{y}(\tau+\ell|t)$ by setting them to the mean of future outputs. In this way the predictors minimize the MSE on output prediction. Using quadratic losses, the ideal control cost function to be minimized w.r.t. $u$ over an horizon of length $T$ is formulated as:
$J_{u}\!$ $=\!\sum_{\ell=0}^{T-1}\|y_{\text{ref}}(\tau+\ell)-\hat{y}(\tau+\ell)\|_{Q_{\ell}}^{2}\!+\!\sum_{\ell=0}^{T-1}\|u_{\text{ref}}(\tau+\ell)-u(\tau+\ell))\|_{R_{\ell}}^{2}$ (5)
@@
where here, and in what follows, $\hat{y}(\tau+\ell)$ is used in place of $\hat{y}(\tau+\ell|\tau)$ to simplify notation. In (5) the outputs can now be also vectors, $Q_{\ell}$ and $R_{\ell}$ are positive semidefinite matrices, and $\|x\|_{M}^{2}:=x^{\top}Mx$ for any matrix $M$. Finally, $u_{\text{ref}}$ are reference inputs that define the penalty on the energy required by the control.
@@
In a more realistic MPC scenario $\mathcal{S}$ is not known and system identification can enter the scene first. Following our previous discussion, we can introduce a parametric model structure $\mathcal{M}(\theta)$ to capture $\mathcal{S}$. All the induced predictors become $\hat{y}(\tau+\ell;\theta)$, function of the unknown vector $\theta$. In turn, this leads to the parametrized control cost $J(u;\theta)$. In this setting the simplest version of the so-called indirect approach to data-driven control, of which the classical identification mentioned above is just an example, first determines $\theta$ from the input-output data $\mathcal{D}$ collected in the previous experiment. This can be done through PEM, i.e. minimizing the one-step-ahead prediction error as in (2), possibly solving the more complex problem (2) in the case of unknown model complexity. Then, the objective $J(u;\theta)$ is minimized by replacing $\theta$ with its estimate (the realization of $\hat{\theta}$).
@@
There are some recurring criticisms of the indirect approach that have emerged in recent years. The first one is that bias is introduced in the control objective [13]. Note however that such issue depends on the indirect approach adopted. For instance, classical system identification is parametric in nature but one could also consider different non parametric and/or Bayesian settings as we will do later on. Furthermore a careful definition of bias is needed. To this regard,
@@
in the Fisherian context where $\theta$ is deterministic, we say that the indirect approach to control is unbiased if, for any input $u$, one has E[J(u;^θ)]=J(u;θ)  θ, with expectation taken w.r.t. the distribution of the estimator $\hat{\theta}$.
@@
Assume that the model structure $\mathcal{M}$ contains $\mathcal{S}$ (so that there exists $\theta$ leading to the optimal predictors). Then, even if $\theta$ is unknown, null bias guarantees that the mean of the cost defined using $\hat{\theta}$ corresponds to the ideal objective (5).
@@
Adopting the classical PEM approach, the unbiasedness property hardly holds because $J$ is a nonlinear function of $\theta$. On the other hand, when e.g. the identification data $\mathcal{D}$ are nonlinearly related to the model parameters, $\hat{\theta}$ is rarely an unbiased estimator of $\theta$ itself [^2] This is by no means in itself a negative thing. We are here concerned with unbiased costs but there exist estimators that introduce a small bias to greatly decrease variance, hence leading to a favourable MSE. Stein also proved in the linear regression setting that, if the dimension of $\theta$ exceeds two, the least squares estimator is not admissible: there are biased estimators that have smaller MSE for any $\theta$, e.g. see [16] and [33, Chapter 1].. The rare exceptions include linear regression problems with linear least squares adopted to estimate $\theta$. For this reason one could try to remove this distortion exploiting the experimental observations in a different way. Indeed, the idea that permeates the direct approach is that the information contained in the data should be maximally exploited by focusing directly on the control cost. So,
@@
the problem of relevance related to the direct approach to data-driven control is to build directly from input-output data an estimator of (ideal) control objectives like (5) having desired statistical properties (like unbiasedness).
@@
4 Separation principle for identification and control